OpenStudy (anonymous):

if cos theta=-5/13 and theta is greater than or equal to 0 and less then or equal to 2pi, determine a) Tan theta b) sec theta

OpenStudy (anonymous):

\[\cos \theta=\frac{ -5 }{ 13 }, 0\le \theta \le2\pi\] find: \[a)\tan \theta\] \[b)\sec\theta\]

OpenStudy (anonymous):

What I think it should be for tan theta tan theta= 12/-5?

OpenStudy (anonymous):

Sec is just 1/ Cos so Sec theta = 13/-5

OpenStudy (anonymous):

is my answer correct for tan theta cuz the book got a different answer

OpenStudy (anonymous):

Tan = sin / cos Since the sin is Opposite/Hypotenuse Sin = 12/13 Cos is Adjacent/Hypotenuse so it is -5/13 12/13 * 13/-5 = -12/5 = 12/-5 so yes you did get the correct answer for tan theta

OpenStudy (anonymous):

the book says its 12/13

OpenStudy (anonymous):

for tan theta

OpenStudy (anonymous):

oh well i guess they messed up.

OpenStudy (anonymous):

Yeah I think they meant to say that Sin theta = 12/13